OpenStudy (anonymous):

Assuming x ≠ 0 and y ≠ 0, what is the quotient of 8x^5y^2+12x^3y^5+20x^2y^7/4xy^2

OpenStudy (zepp):

Let's put it in LaTex :"( \(\huge \frac{8x^5y^2+12x^3y^5+20x^2y^7}{4xy^2}\)

OpenStudy (zepp):

Then break it up \(\huge \frac{8x^5y^2}{4xy^2}+\frac{12x^3y^5}{4xy^2}+\frac{20x^2y^7}{4xy^2}\)

OpenStudy (zepp):

Then MASS SIMPLIFY \(\huge \frac{4x^4}{1}+\frac{3x^2y^3}{1}+\frac{5xy^5}{1}\)

OpenStudy (anonymous):

wow ok that makes ALOT MORE sense thank you